OpenStudy (anonymous):

2x+5=7 find the value of x

OpenStudy (lgbasallote):

2x + 5 = 7 subtract 5 from both sides 2x = 7 - 5 2x = 2 divide both sides by2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

thnks

OpenStudy (lgbasallote):

welcome :D